OpenStudy (lillian_a):

Which of the following functions is not continuous for all real numbers x ? A) f(x) = x^1/3 B) f(x) = 2/(x + 1)^4 C) f(x) = |x + 1| D) f(x) = √(1 + e^x) E) f(x) = x − 3/(x^2 + 9) How do i figure out if a function is continuous at all real numbers?

imqwerty (imqwerty):

if a function is continuous it is defined for every value of x we can also say that we can draw the graph of a continuous function without lifting our pen for example -this function is continuous

OpenStudy (lillian_a):

what method would i use to find that is has continuity

imqwerty (imqwerty):

are you allowed to use graphing calculators?

OpenStudy (lillian_a):

No

OpenStudy (solomonzelman):

If you see that a function is undefined at a particular x-vaue, then the function is not continous.

OpenStudy (solomonzelman):

For example, \(\large\color{black}{ \displaystyle y=\frac{1}{x-3} }\) is not continous, because the function is undefined at x=3. (you end up dividing by 0)

OpenStudy (lillian_a):

how did you get that?

OpenStudy (solomonzelman):

That is an example, not an answer...

OpenStudy (lillian_a):

I know

OpenStudy (solomonzelman):

Can you divide by 0?

OpenStudy (lillian_a):

No...

OpenStudy (solomonzelman):

Yes, you can NOT! good! When, the function: \(\large\color{black}{ \displaystyle y=\frac{1}{x-3} }\) is evaluated at x=3, you get: \(\large\color{black}{ \displaystyle y=\frac{1}{\color{red}{3}-3} =\frac{1}{0}\color{blue}{=\rm UNDEFINED} }\)

OpenStudy (lillian_a):

So if you plug zero into the equation and it's undefined then it not continuous?

OpenStudy (solomonzelman):

So when x=3, your output (y) is undefined. Therefore, the function is not continous because it has a hole at x=3.

OpenStudy (solomonzelman):

Wel, if for some vaues of x, let's say x=c, you get an undefined output, THEN this means that the function is not continous because it has a gap at x=c.

OpenStudy (lillian_a):

But I'm not given a point...

OpenStudy (solomonzelman):

you know that in a case of, (lets make a new function up) y=1/(x+5) Wha happenes when you plug in x=-5 into this function?

OpenStudy (lillian_a):

you get undefined

OpenStudy (solomonzelman):

yes, exacty that means that you can't have a point for x=-5, right?

OpenStudy (solomonzelman):

in other words, there is nothing there on the vertical of x=-5.

OpenStudy (solomonzelman):

And that means you have a gap for x=-5. And this means the function is not continous (because of the hole at x=-5)

OpenStudy (lillian_a):

but i just trying to figure out which ONE is not continuous

OpenStudy (solomonzelman):

Ok, analyze each function in the answer-choices. You might have an "undefined" output, for a coupe of reasons. 1) Negative in the square root (negative in the 4th roots, or in any even root) 2) When you end up dividing by 0

OpenStudy (solomonzelman):

Note that negative inside the cube, fifth or any ODD rood, would still be defined

OpenStudy (solomonzelman):

lets do option 1 together

OpenStudy (solomonzelman):

A) f(x) = x^1/3 B) f(x) = 2/(x + 1)^4 C) f(x) = |x + 1| D) f(x) = √(1 + e^x) E) f(x) = x − 3/(x^2 + 9) I wil post a options for reference

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle x^{1/3}}\)

OpenStudy (solomonzelman):

Will you end up "dividng by zero" if this function does not even have a denominator?? Is this function even undefined if you pug lets say -8 or -1 for x, or if you plug anything else into this function?

OpenStudy (lillian_a):

no

OpenStudy (solomonzelman):

No to which question(s), the first, the second or both?

OpenStudy (lillian_a):

both

OpenStudy (solomonzelman):

very good, so that means that this function is _____ ?

OpenStudy (lillian_a):

the function is continuous

OpenStudy (solomonzelman):

Yes, very good, answer choice A is continous

OpenStudy (solomonzelman):

Which letter do you want to do next?

OpenStudy (lillian_a):

c

OpenStudy (solomonzelman):

\(\large\color{#000000 }{ \displaystyle f(x)=\left|x+1\right| }\)

OpenStudy (solomonzelman):

We, that is just an absolute value (bars denote an absolute value)

OpenStudy (solomonzelman):

Do you think there is a number that you can plug into the function for x, to make the function undefined ?